++++ INAF LBC+LUCIFER queue observing, 2011 January 30 - 31

Observer: K. Boutsia, R. Speziali
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LUCIFER

Summary

Mostly clear at the beginning of the night but thick clouds are approaching (visible now at the horizon)
Clouds are passing by throughout the night
Dome remained closed for most of the night due to clouds - high wind - high humidity
During operational time we had problems with the calibration unit - we found it switched off and it had to be restarted manually

Data Summary

Project Objects Filter ExpTime Notes

Turatto SN2011B J, H, Ks 12 s + 12 s + 24 s  
Turatto SN2011B 200HK , orderSep 2000sec arc+flat+tell
Turatto SN2010al J, H 640 sec+ 504sec clouds passing by - H imaging interrupted - dome closed

Details

Begin night with LUCIFER

12:50 - Sky flats
  • Ks flat 0002-0011 high (16000counts) 0012-0021 medium (10000) 0022-0031 low (5000) - here the type keyword is DARK !!!
  • H flats 0032-0041 high (16000) 0052-0061 low (5000) - the keyword has been correctly set to FLAT
  • J flats 0042-0051 high (20000-16000) 0062-0081 med and low (7000 -4000)

01:12 - Pointing correction near the SN2011B field (Turatto)

01:23 - Image Acquisition script for SN2011B (Turatto) 0082-0083
  • Sky is still too bright - no guide star found
  • we wait and in the meantime we correct pointing again and collimate on Perssons star

01:36 - repeat image acquisition SN2911B 0084-0085
  • seeing on the guide star is around 1arcsec

01:40 - Science imaging scripts for 2011B
  • J imaging 0086-0089
  • H imaging 0090-0093
  • Ks imaging 0094-0097

01:48 - Spec Acquisition script 2011B 0098-0100

01:57 - Spec Science script 2011B 0101-0108
  • one extra image to see if object in the slit 0109-0111 no good because grating still in...

02:36 - CALIB UNIT IN
  • problems with the calibration unit - 10minutes lost
  • the calib unit had set itself off and had to be manually restarted
  • ARC 0112-0119 here the PI is KBoutsia because the archive information was missing
  • FLAT 0120-0134

02:55 - Acquisition of the telluric for SN2011B 0135-0138

03:06 - Telluric Science acquisition 0139-0146
  • we have changed the offsets so that the first spectrum does not end in the saturated part of the detector (0 50 / 0 -100)
  • too large offset - lost guide star
  • repeat with smaller offsets (0 30 / 0 30)

03:13 - Slew to SN2010al (Turatto) for image acquisition 0147-0152

03:22 - Science imaging for SN2010al
  • J 0153-0158
  • during the 4th offset is getting quite cloudy - the guide star is blurry
  • the script failed during the 7th offset because could not find the guiding star

03:36 - too cloudy - we can't see the guide star and we wait
  • the wind is picking up (18m/sec)

03:44 - we can see the guide star again - continue science imaging for SN2010al
  • J imaging 0159-0163 we omitted the first 3 offsets because they should be fine - 5 offsets this time to finish the OB
  • H imaging 0164-0172 clouds going through - 8th and 9th offset not much to see - at 10th offset it quit

04:12 - we wait for the guide star to appear again

04:30 - clouds are getting thicker - DOME closed

07:40 - Opening Dome

07:45 - Pointing Correction
  • the tertiary is giving errors - probably is moving too much because of the wind

08:02 - Spec acquisition for SDSSJ1048+4637 (Gallerani)
  • we can not see the guide star
  • we try with alternative guide star but we still cannot see anything

08:08 - Clouds are building up again and the wind is getting stronger - closing dome again

13:08 - Dark frames - End of the night
  • the dark frames tonight were taken with the blinds off and are not reliable (frames - luci_20110131_Dark_???.fits)
  • dark frames for tonight's images were obtained at the beginning of the following night and can be found in the directory /data/luci/20110201/

-- VincenzoTesta - 31 Jan 2011
Topic revision: r5 - 01 Feb 2011, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback